a boat goes 350 miles downstream in the same time go 250 miles upstream . The speed of the current is 5 miles per hour. Find the speed of the boat in still water. I know the answer is 45 mph but i want to know how they got it.
Answer by scott8148(6628)

t = d / r ___ the times are equal

350 / (b + 5) = 250 / (b - 5)

"cross" multiplying ___ 350b - 1750 = 250b + 1250

100b = 3000


the answer you "know", appears to be incorrect